the search coil has 200 turns and a cross sectional area of 3.5 x 10-5 m2. the search coil is placed at x = 0.070 m. show that the magnetic flux linkage through the search coil is about 5x 10-4Wb


1
Expert's answer
2022-03-28T14:25:47-0400

Solenoid

Magnetic field

"B=\frac{\mu_0 ni}{2x}"

"\frac{\phi}{A}=\frac{4\times3.14\times200\times
 i}{2\times0.070}"

"\frac{5\times10^{-4}}{3.5\times10^{-5}}=\frac{4\times3.14\times200\times
 i}{2\times0.070}"

"I=7.96\times10^{-4}A"

"B=\frac{\mu_0n I}{2r}"


"B=\frac{4\times3.14\times10^{-7}\times200\times7.96\times10^{-4}}{2\times0.070}=1.42\times10^{-6}T"

Magnetic flux

"\phi=BA"

"\phi=5\times10^{-4}Wb"